OpenStudy (anonymous):

Help Please! Simplifying Radical Expressions: (sqrt)35x^5y^2

OpenStudy (unklerhaukus):

sqrt(35x^5y^2) ?

OpenStudy (anonymous):

Mhmm.

OpenStudy (unklerhaukus):

is that what you meant? or sqrt(35)x^5y^2

OpenStudy (anonymous):

\[\sqrt{35x^5y^2}\]

OpenStudy (unklerhaukus):

ah ok , do you know that these are equilavalent? \[\sqrt (x)=(x)^{1/2}\]

OpenStudy (anonymous):

Yes sir.

OpenStudy (unklerhaukus):

so you can write your expression something like \[(ab^nc^m)^p\] now you just have to multiply through that power \[=a^pb^{n\times p}c^{m\times p}\]

OpenStudy (anonymous):

Hmm. I've never seen that before.

OpenStudy (anonymous):

I made a mistake. \[\sqrt{32x^5y^2}\]

OpenStudy (unklerhaukus):

so \(a\) is like 32

OpenStudy (anonymous):

Okay. I follow you.

OpenStudy (unklerhaukus):

what do you get when you change the radical sign in the expression to power (index)

OpenStudy (anonymous):

Can you give me an example of how this works?

OpenStudy (unklerhaukus):

this is a similar example \[\sqrt[3]{8w^6v^2}=(8w^6v^2)^{1/3}=8^{1/3}w^{(6\times1/3)}v^{(2\times1/3)}=2w^2v^{2/3}\]

OpenStudy (unklerhaukus):

first step was to change the radical (this time cube root) to index (power) form, then i multiplied the power through the brackets and finally simplified the indices in your question the number term wont simplify as nicely as my example did

OpenStudy (unklerhaukus):

@Liv_16 still here?